OpenStudy (anonymous):

what equation is parallel to -3+4x=y+2x and has a y-intercept of 9 ?

OpenStudy (anonymous):

-3+4x=y+2x is equivalent to y= 2x + 3 so any function 2x + b is parallel

OpenStudy (anonymous):

i meant 2x -3 !!! sorry! ^^"

OpenStudy (anonymous):

thank you .

OpenStudy (anonymous):

y-intercept: y= -3 (you see that if x=0 since 2*0 -3 = -3 )

OpenStudy (kainui):

So when you look at y=mx+b you know that y and x are variables that you have no control over, so you shouldn't worry about them so much. What you need to worry about is what you pick for m and b. So what's m? What's b? m is the slope and b is the y-intercept. So if you change m and keep b the same, your graph looks like this: If you change b while keeping m the same, it looks like this:  So there's some food for thought. Play around with a graphing calculator just for about a minute trying different slopes and intercepts and you will understand and look like a god to your classmate friends. Promise.